Select the approximate values of x that are solutions to f(x) = 0, where
f(x) = -9^2 + 2x + 8.
help please!

Answer by rfer(16322)

0=81+2x+9
0=90+2x
-90=2x
-45=x
--------------
0=81+2*-45+9
0=90-90